OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = three halvesx + 4 and passes through the point (-6, 3). y = three halvesx - 1 y = three halvesx + 3 @gorv

OpenStudy (gorv):

y=3x/4+4

OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 3x - y = -3 and passes through the point (4, -2). y = -one thirdx - 6 y = -one thirdx - 14 y = 3x - 6 y = 3x - 14

OpenStudy (anonymous):

oops sorry, meant that one.

OpenStudy (gorv):

product of slope of perpendicular lines=-1 let slope of perpendicular line =m

OpenStudy (gorv):

convert it in standard form y=mx+c

OpenStudy (gorv):

add 3 on both side wat u got ??

OpenStudy (anonymous):

3x & 0.

OpenStudy (gorv):

3x-y=0

OpenStudy (gorv):

nope u have to add on both side

OpenStudy (gorv):

3x-y+3=0 this will we get

OpenStudy (gorv):

now add y on both side

OpenStudy (anonymous):

I just got so confused lol.

OpenStudy (anonymous):

Running out of time on the test too.. :/

OpenStudy (gorv):

3x+3=y slope =3

OpenStudy (gorv):

y=3x+c equation of parallel line

OpenStudy (gorv):

itzz passing through (4,-2) -2=3*4+c -2=12+c c=-14

OpenStudy (gorv):

y=3x-14

OpenStudy (anonymous):

Okay, thank you. :)

OpenStudy (gorv):

welcome